OpenStudy (anonymous):

an aqueous BeCl2 solution has a mass percentage of 8% and a density of 1.10g/mL. A person takes 10mL of the solution and adds water till the total volume is 500mL, what is the molarity of BeCl2?

OpenStudy (anonymous):

@maheshmeghwal9 something we can work on together? :)

OpenStudy (maheshmeghwal9):

Wt did u do upto this?

OpenStudy (maheshmeghwal9):

An aqueous solution of BeCL2 = water + BeCL2.

OpenStudy (maheshmeghwal9):

so now the question is more easy to think upon:)

OpenStudy (anonymous):

yes, so i know that: \[\rho=1.10g/mL and that \rho=m/v\] \[molarity= mass solute/L solution (M)\] mass % = 8% =(mass solute/mass solution) x100 V1=10mL and V2=500mL

OpenStudy (maheshmeghwal9):

I think u should let BeCl2's mass be 'x'. do u agree?

OpenStudy (anonymous):

so i got the mass of BeCl2 by solving the density equation for mass: m=vp = 1.10 x 500?

OpenStudy (vincent-lyon.fr):

You can solve the problem in that way: Each value you obtain goes in the next line. - 10 mL of solution have a mass of.........................................g - ....... g of solution have a mass of BeCl2 of............................g - ........g of BeCl2 represent an amount of BeCl2 of...................mol. - ....... mol of BeCl2 in 500mL make a concentration of.............mol/L

OpenStudy (anonymous):

@Vincent-Lyon.Fr , this is how far I got with the explanation: 10mL of BeCl2 solution has a mass of 11g. I don't understand what you mean by the 2nd part.

OpenStudy (vincent-lyon.fr):

Next step is : 11 g of solution contain how many g of BeCl2 ? Remember mass percentage is 8%

OpenStudy (anonymous):

oo okay, then it contains 0.88gBeCl2! thank you!!^^